The results of a state-wide examination of the reading skills of sixth-grade students were

normally distributed, with a mean score of 104 and a standard deviation of 16. The

students in the top 10% are to receive an award, and those in the bottom 14% will be

required to take a special reading class.

a. What score does a student need in order to receive an award?

b. What is the cut-off score that will be used to determine whether a student will be

required to take the special reading class?


1
Expert's answer
2021-12-15T12:39:11-0500

Consider that the mean of the grades is 104 and the standard deviation is 16. The students in the top 10% are to receive an award, and those in the bottom 14% will be required to take a special reading class.

a. The following figure shows a standard normal distribution with 10% of the data to the right of some unknown z-score and 40% of the data to the left of the z-score but to the right of the mean of 0.



Hence, the unknown z-score is right of 0, thus

"z_x = 1.28"

Let x represents the score for which a person receive an award, then x is related to the z-score by the formula

"z_x = \frac{x -\bar{x}}{s} \\

1.28 = \frac{x-104}{16} \\

20.48 = x-104 \\

x = 124.48"

Hence, a student needs a score of 124.48 in order to receive an award.

b. The following figure shows a standard normal distribution with 14% of the data to the left of some unknown z-score and 36% of the data to the right of the z-score but to the left of the mean of 0.



Hence, the unknown z-score is to the left of 0, thus

"z_x = -1.08"

Let x represents the score for which a person receive an award, then x is related to the z-score by the formula

"z_x = \frac{x -\bar{x}}{s} \\

-1.08 = \frac{x-104}{16} \\

-17.28 = x -104 \\

x = 86.72"

Hence, the cut-off score to determine the requirement of special reading class is 86.72